OpenStudy (anonymous):

I have a question on continuous compounding. The problem asks me to calculate n for the formula A=Aok^n when A (the amount) doubles. I know I have to do logs on both sides which gives me 2=2k^n but I don't know where to go from there.

OpenStudy (amistre64):

2P = Pe^(rt) 2 = e^(rt), when rt = ln(2) solving for t ... t = ln(2)/r

OpenStudy (amistre64):

i take it Ao is the 'P'rincipal amount

OpenStudy (amistre64):

2A = Ao k^n 2A/Ao = k^n log (2A/Ao) = n log(k) log (2A/Ao) ---------- = n log(k)

OpenStudy (amistre64):

i spose 2A is a bad thought ... let A = 2Ao, makes more sense, so just log(2) on top